Lhyn:

Find the number of edges and the number of vertices of a regular polyhedron whose number of faces, each of which a triangle, is 20

Vocaloid:

@tranquility would you mind taking a look at this? I know Euler's formula exists for F + V = E + 2 but I'm not sure how to proceed with only the face count

Vocaloid:

Plagiarized answers are not acceptable here.

Vocaloid:

Alright, now that I’m looking at this with a fresh set of eyes and I’ve had some opportunity to research more about polyhedrons : On this polyhedron, each edge is shared by a unique set of two triangles. Therefore, taking 20 faces * 3 edges per face you get 60 possible edges, however, because each edge is shared between two triangles there are only 60/2 = 30 edges. Now that you have the face count and the edge count you can go back to Euler’s formula and solve for V to get vertices. Notice how I am alluding to the stack exchange answer without copy-pasting word for word.